# VideoPrism: A Foundational Visual Encoder for Video Understanding

### v1

@@ Abstract
We introduce VideoPrism, a general-purpose video encoder that tackles diverse video understanding tasks with a single frozen model. We pretrain VideoPrism on a heterogeneous corpus containing 36M high-quality video-caption pairs and 582M video clips with noisy parallel text (e.g., ASR transcripts). The pretraining approach improves upon masked autoencoding by global-local distillation of semantic video embeddings and a token shuffling scheme, enabling VideoPrism to focus primarily on the video modality while leveraging the invaluable text associated with videos. We extensively test VideoPrism on four broad groups of video understanding tasks, from web video question answering to CV for science, achieving state-of-the-art performance on 30 out of 33 video understanding benchmarks.
@@ 1 Introduction
Contributions.
VideoPrism is a state-of-the-art, general-purpose video encoder. We advocate for a scalable strategy for collecting pretraining videos, combining manually captioned videos with those containing noisy textual descriptions. We design a unique two-stage pretraining approach tailored to this hybrid data, leveraging video-language contrastive learning to harvest semantics, followed by improved masked video modeling with global-local distillation and token shuffling. Finally, we present a comprehensive evaluation on four broad categories of understanding tasks across 33 diverse benchmarks, including videos from the web, scripted performances, and scientific experiments. Results demonstrate that VideoPrism significantly outperforms existing ViFMs on 30 benchmarks (Figure 2). Importantly, no single baseline model consistently achieves second-best performance, indicating VideoPrism’s robust generalizability.
@@ 2.1 Pretraining data
Our pretraining data consists of 36M clips (sampled from 36M videos) with high-quality manually labelled captions and 582M clips (from 275M videos) with noisy parallel text, as summarized in Table 1. The 36M high-quality video-caption pairs in Anonymous-Corpus #1 are the largest of its kind for ViFMs, to our knowledge, but they are still an order of magnitude smaller than the image-language data used to fuel image FMs (Radford et al., 2021; Yu et al., 2022). Hence, we also collect large-scale video-text data whose noisy text is generated through ASR, metadata, and large multimodal models (Wang et al., 2023e; Zhao et al., 2024), etc. This subset of videos correspond to the rows from WTS-70M to Anonymous-Corpus #3 in Table 1, and we provide more details in Appendix A.
Importantly, unlike previous works (Tong et al., 2022; Wang et al., 2022c; Li et al., 2023b; Wang et al., 2023b), we do not incorporate any training sets from the evaluation benchmarks, e.g., Kinetics (Kay et al., 2017), for either pretraining or post-pretraining. This conscious choice avoids overly tuning our model towards certain evaluation benchmarks. Moreover, we carefully de-duplicate the pretraining corpus against the videos in the evaluation benchmarks used in this paper to ensure that there is no data leakage.
@@ 2.2 Model architecture
Figure 3: Illustration of our two-stage pretraining. Stage 1 trains video and text encoders with contrastive loss on video-text pairs, supplying semantic video embeddings to the next stage. Stage 2 continues to train the video encoder using improved masked autoencoding on video-only clips. The encoder uses unmasked 3D video patches to predict a global semantic embedding of the whole video and token-wise embeddings. Decoder 2 processes shuffled tokens with positional embedding, while Decoder 1 has no positional embedding.
@@ 2.3 Training algorithm
Our goal is to leverage both video-text pairs and the video-only data curated in Section 2.1 to train VideoPrism scalably, so as to make VideoPrism a foundational video encoder capable of capturing both appearance and motion semantics from videos. We highlight the video-only modality rather than solely relying on video-text because the text in our large-scale pretraining corpus is very noisy for some videos. As shown in Figure 3, the training pipeline of VideoPrism includes two stages: video-text contrastive training and masked video modeling.
@@ 2.3.2 Stage 2: Masked video modeling
In this stage, we continue to train the video encoder on video-only data using improved masked video modeling. These improvements consist of (1) a novel token shuffling scheme to prevent decoding shortcuts and (2) global and token-wise distillation losses to effectively leverage the knowledge acquired in the first stage. As illustrated in Figure 3, the second-stage (student) model learns to predict the first-stage (teacher) model’s embeddings of all tokens based on a masked video. The encoder-decoder Transformers are decoupled following He et al. (2022)’s design.
@@
As we effectively initialize the second-stage model from the first stage, one issue is that the model may create a shortcut for the decoder to copy and paste the unmasked tokens while predicting only the masked ones, making it an easier task to solve than predicting all tokens. To address this issue, we randomly shuffle the token sequence output by the encoder before feeding it to the decoder, and the decoder adds positional embeddings to this sequence after the shuffling. Note that this shuffling operation avoids the copy-and-paste shortcut of unmasked tokens that the decoder can potentially explore. One can also view it akin to Jigsaw puzzles (Noroozi & Favaro, 2016) for the unmasked tokens while predicting the masked ones.
@@
3 Experiments
We evaluate VideoPrism on a wide spectrum of video-centric understanding tasks to demonstrate its capability and generalizability. We group the tasks into four categories: (1) general video-only understanding, including classification and spatiotemporal localization (Section 3.1), (2) zero-shot video-text retrieval (Section 3.2), (3) zero-shot video captioning and QA (Section 3.3), and (4) CV for science (Section 3.4). For all experiments in the main paper, we freeze VideoPrism as a video encoder and only train task-specific components for the tasks in groups (1), (2), and (4) and some adaptation layers connecting VideoPrism to an LLM for (3). In the appendices, we report more results of end-to-end and adapter finetuning. Note that our evaluation strategy, freezing the visual encoder, aligns with prior works (He et al., 2022; Singh et al., 2022; Yuan et al., 2023) and is almost a go-to choice for building VideoLLMs (Tang et al., 2023). It is especially needed for videos because finetuning a ViFM is prohibitively expensive, while a frozen ViFM allows one to amortize the cost of video encoding across multiple tasks.
@@ 3.1 Classification and spatiotemporal localization
We compare VideoPrism with state-of-the-art FMs on a video-only understanding benchmark: VideoGLUE (Yuan et al., 2023). By design, VideoGLUE evaluates FMs through four adaptation methods over eight hallmark datasets, representing appearance-focused action recognition (VC (A)), motion-rich action recognition (VC (M)), multi-label video classification (VC (ML)), temporal action localization (TAL), and spatiotemporal action localization (STAL). Moreover, this benchmark introduces a VideoGLUE score (VGS), considering the tradeoff between adaptation costs and performance, to provide a holistic view of FMs’ capabilities on the video-only understanding tasks. We present the frozen-backbone evaluation results in the main paper and leave the rest to Appendix E. We employ an MAP head (Yuan et al., 2023) in action recognition (MAP probing) and spatiotemporal localization and use G-TAD (Xu et al., 2020) for temporal localization (see Section E.1 for details).
@@ 3.2 Zero-shot video-text retrieval and classification
Table 3: Results of zero-shot video-text retrieval. We report the Recall@1 (R@1) and R@5 for all the benchmarks.
Methods MSRVTT VATEX ActivityNet
@@
VideoPrism-B 37.0 61.5 67.7 87.5 57.7 88.5 76.2 93.7 49.6 76.7 47.9 75.0
($\downarrow$5.6) ($\downarrow$2.9) ($\uparrow$3.0) ($\uparrow$2.3) ($\uparrow$4.5) ($\uparrow$5.2) ($\uparrow$2.6) ($\uparrow$0.5) ($\uparrow$6.8) ($\uparrow$7.1) ($\uparrow$7.2) ($\uparrow$7.4)
VideoPrism-g 39.7 63.7 71.0 90.0 62.5 91.0 77.1 95.6 52.7 79.4 50.3 77.1
($\downarrow$2.9) ($\downarrow$0.7) ($\uparrow$6.3) ($\uparrow$4.8) ($\uparrow$9.3) ($\uparrow$7.7) ($\uparrow$3.5) ($\uparrow$2.4) ($\uparrow$9.9) ($\uparrow$9.8) ($\uparrow$9.6) ($\uparrow$9.5)
@@
We evaluate VideoPrism’s zero-shot video-text retrieval performance on three benchmarks: MSRVTT (Xu et al., 2016), VATEX (Wang et al., 2019), and ActivityNet (Caba Heilbron et al., 2015). For zero-shot video classification tasks, we experiment with Kinetics-400 (Kay et al., 2017), Charades (Sigurdsson et al., 2016), SSv2-Temporal and SSv2-Events (Sevilla-Lara et al., 2021; Bagad et al., 2023), and the ATP-Hard subset of NExT-QA (Buch et al., 2022). SSv2 and NExT-QA (ATP-Hard) focus on motion and temporal reasoning, respectively. Moreover, we adapt Charades-STA (Gao et al., 2017) to the zero-shot classification scenario by reformulating each of its samples in the test set into a multi-choice retrieval problem (see Section F.2 for more details). We report results following the standard evaluation metric for each benchmark.
@@
Tables 3 and 4(e) summarize the results of video-text retrieval and video classification, respectively. VideoPrism sets the new state of the art on most benchmarks, and the gains over the prior arts are exceptionally substantial on the challenging datasets (e.g., $9.5\%$ on ActivityNet, $4.4\%$ on SSv2-Events, and $6.6$ mAP on Charades). Most results from our base-scale VideoPrism-B are actually better than those of existing larger scale models. Additionally, VideoPrism is on par with or better than the models pretrained with in-domain data and extra modalities (e.g., audios) in Table 4(e). These improvements in zero-shot retrieval and classification tasks present VideoPrism’s strong generalization capabilities.
@@ 3.3 Zero-shot video captioning and QA
We evaluate the model in the zero-shot setting on the test splits of a suite of standard video captioning datasets including MSRVTT (Xu et al., 2017), VATEX (Wang et al., 2019), and YouCook2 (Zhou et al., 2018), and video QA benchmarks including MSRVTT-QA (Xu et al., 2017), MSVD-QA (Xu et al., 2017), and NExT-QA (Xiao et al., 2021). For video QA, where it is imperative to match the length and style of the model’s answers with groundtruths, we adopt the zero-shot approach of Flamingo (Alayrac et al., 2022) and use two-shot text-only prompts from the training set of the downstream task. Additionally, for MSRVTT-QA and MSVD-QA, we experiment with the closed-vocabulary evaluation configuration (Li et al., 2022; Yang et al., 2022). In this setting, we let the model score candidate answers according to their log-likelihoods and return the top one.
@@ 3.6 Limitations
One limitation of our approach is that we leverage a video corpus with noisy text as part of pretraining. This noisy text is potentially incomplete and biased, which could impact model performance. Moreover, long video understanding remains a challenge, since our current focus is on short video clips from which we sample $16$ frames as input to VideoPrism. Future work in this direction could leverage our encoder as part of a long video understanding system. Finally, while we advocate for the frozen-backbone evaluation, we acknowledge that there are scenarios that benefit more from end-to-end finetuning and parameter-efficient adaptation. Despite these limitations, results demonstrate the potential impact of VideoPrism has on a range of real-world video understanding tasks.
@@ 4 Related work
Foundation models (FMs) (Bommasani et al., 2021) have demonstrated tremendous promise with early work in LLMs (Devlin et al., 2019; Brown et al., 2020). Some ViFMs are built around LLMs (Wang et al., 2022d; Li et al., 2023a; Zhang et al., 2023; Chen et al., 2023a), analyzing videos by feeding associated text to LLMs, such as ASR transcripts and machine-generated captions. In contrast, VideoPrism takes a video-centric view, and we aim to tackle a broader range of video understanding tasks.
ViFMs.
Most recent FMs in CV focus on images (Radford et al., 2021; Yuan et al., 2021; Jia et al., 2021; Yu et al., 2022; Alayrac et al., 2022; Yan et al., 2022; Wang et al., 2022a; Chen et al., 2023c; Xu et al., 2023; Girdhar et al., 2023; Zhu et al., 2024). Their pretraining data contains no or only a small portion of videos, and the model architectures and learning methods are for images by design. While these FMs can accept video frames as input, they fall short on motion and temporal modeling (Yuan et al., 2023). Our work directly addresses this gap by developing a video encoder.
@@
Pretraining strategy.
Our pretraining integrates vision-language contrastive learning (Radford et al., 2021; Xu et al., 2021; Bain et al., 2022) and masked data modeling (Devlin et al., 2019; He et al., 2022). The former has led to strong late-fusion models like CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), CoCa (Yu et al., 2022), and the latter is proven effective to learn from single-modality data like language (Devlin et al., 2019; Anil et al., 2023), audio (Borsos et al., 2023), images (He et al., 2022; Wang et al., 2023d; Oquab et al., 2023), and videos (Tong et al., 2022; Wang et al., 2023b). While EVA (Fang et al., 2022, 2023) and UMT (Li et al., 2023b) transfer indirect semantics from CLIP (Radford et al., 2021) to masked modeling, we learn video-native semantics. We also introduce global distillation and token shuffling to the masked video modeling to orchestrate both appearance and motion cues.

### v3

@@ Abstract
We introduce VideoPrism, a general-purpose video encoder that tackles diverse video understanding tasks with a single frozen model. We pretrain VideoPrism on a heterogeneous corpus containing 36M high-quality video-caption pairs and 582M video clips with noisy parallel text (e.g., ASR transcripts). The pretraining approach improves upon masked autoencoding by global-local distillation of semantic video embeddings and a token shuffling scheme, enabling VideoPrism to focus primarily on the video modality while leveraging the invaluable text associated with videos. We extensively test VideoPrism on four broad groups of video understanding tasks, from web video question answering to CV for science, achieving state-of-the-art performance on 31 out of 33 video understanding benchmarks. Our models are released at https://github.com/google-deepmind/videoprism.
@@ 1 Introduction
Contributions.
VideoPrism is a state-of-the-art, general-purpose video encoder. We advocate for a scalable strategy for collecting pretraining videos, combining manually captioned videos with those containing noisy textual descriptions. We design a unique two-stage pretraining approach tailored to this hybrid data, leveraging video-language contrastive learning to harvest semantics, followed by improved masked video modeling with global-local distillation and token shuffling. Finally, we present a comprehensive evaluation on four broad categories of understanding tasks across 33 diverse benchmarks, including videos from the web, scripted performances, and scientific experiments. Results demonstrate that VideoPrism significantly outperforms existing ViFMs on 31 benchmarks (Figure 2). Importantly, no single baseline model consistently achieves second-best performance, indicating VideoPrism’s robust generalizability.
@@ 2.1 Pretraining data
Our pretraining data consists of 36M clips (sampled from 36M videos) with high-quality manually labelled captions and 582M clips (from 275M videos) with noisy parallel text, as summarized in Table 1. The 36M high-quality video-caption pairs in Anonymous-Corpus #1 are the largest of its kind for ViFMs, to our knowledge, but they are still an order of magnitude smaller than the image-language data used to fuel image FMs (Radford et al., 2021; Yu et al., 2022). Hence, we also collect large-scale video-text data whose noisy text is generated through ASR, metadata, and large multimodal models (Wang et al., 2023e; Zhao et al., 2024), etc. This subset of videos corresponds to the rows from WTS-70M to Anonymous-Corpus #3 in Table 1, and we provide more details in Appendix A.
Importantly, unlike previous works (Tong et al., 2022; Wang et al., 2022c; Li et al., 2023b; Wang et al., 2023b), we do not incorporate any training sets from the evaluation benchmarks, e.g., Kinetics (Kay et al., 2017), for either pretraining or post-pretraining. This conscious choice avoids overly tuning our model towards certain evaluation benchmarks. Moreover, we carefully de-duplicate the pretraining corpus against the videos in all the 33 evaluation benchmarks used in this paper to ensure that there is no data leakage.
@@ 2.2 Model architecture
Figure 3: Illustration of our two-stage pretraining. Stage 1 trains video and text encoders with contrastive loss on video-text pairs, supplying semantic video embeddings to the next stage. Stage 2 continues to train the video encoder, now called VideoPrism, using improved masked autoencoding on video-only clips. The frozen Stage-1 encoder uses unmasked 3D video patches to produce a global semantic embedding of the whole video and token-wise embeddings. Decoder 2 processes shuffled tokens with positional embedding, while Decoder 1 has no positional embedding.
@@ 2.3 Training algorithm
Our goal is to leverage both video-text pairs and the video-only data curated in Section 2.1 to train VideoPrism scalably, so as to make VideoPrism a foundational video encoder capable of capturing both appearance and motion semantics from videos. We highlight the video-only modality rather than solely relying on video-text because the text in our large-scale pretraining corpus is very noisy for a majority of the videos. As shown in Figure 3, the training pipeline of VideoPrism consists of two stages: video-text contrastive training and masked video modeling.
@@ 2.3.2 Stage 2: Masked video modeling
In this stage, we continue to train the video encoder on video-only data using improved masked video modeling. These improvements include (1) a novel token shuffling scheme to prevent decoding shortcuts and (2) global and token-wise distillation losses to effectively leverage the knowledge acquired in the first stage. As illustrated in Figure 3, the second-stage (student) model learns to predict the first-stage (teacher) model’s embeddings of all tokens based on a masked video. The encoder-decoder Transformers are decoupled following He et al. (2022)’s design.
@@
As we effectively initialize the second-stage model from the first stage, one issue is that the model may create a shortcut for the decoder to copy and paste the unmasked tokens while predicting only the masked ones, making it an easier task to solve than predicting all tokens. To address this issue, we randomly shuffle the token sequence output by the encoder before feeding it to the decoder, and the decoder adds positional embeddings to this sequence after the shuffling. Note that this shuffling operation avoids the copy-and-paste shortcut of unmasked tokens that the decoder can potentially explore. One can also view it akin to Jigsaw puzzles (Noroozi & Favaro, 2016) which the decoder tries to solve for the unmasked tokens while it predicts the masked ones.
@@
3 Experiments
We evaluate VideoPrism on a wide spectrum of video-centric understanding tasks to demonstrate its capability and generalizability. We group the tasks into four categories: (1) general video-only understanding, including classification and spatiotemporal localization (Section 3.1), (2) zero-shot video-text retrieval (Section 3.2), (3) zero-shot video captioning and QA (Section 3.3), and (4) CV for science (Section 3.4). For all experiments in the main paper, we freeze VideoPrism as a video encoder and only train task-specific components for the tasks in groups (1), (2), and (4) and some adaptation layers connecting VideoPrism to an LLM for (3). In the appendices, we report more results of end-to-end and adapter finetuning. Note that our evaluation strategy, freezing the visual encoder, aligns with prior works (He et al., 2022; Singh et al., 2022; Yuan et al., 2023) and is almost a go-to choice for building VideoLLMs (Tang et al., 2023). It is especially needed for videos because finetuning a ViFM is prohibitively expensive, while a frozen ViFM allows one to amortize the cost of video encoding across multiple tasks. All results in the main text are produced using the same frozen VideoPrism-B/g checkpoint, corresponding to the base/giant model.
@@ 3.1 Classification and spatiotemporal localization
We compare VideoPrism with state-of-the-art FMs on a video-only understanding benchmark: VideoGLUE (Yuan et al., 2023). By design, VideoGLUE evaluates FMs through four adaptation methods over eight hallmark datasets, representing appearance-focused action recognition (VC (A)), motion-rich action recognition (VC (M)), multi-label video classification (VC (ML)), temporal action localization (TAL), and spatiotemporal action localization (STAL). This benchmark introduces a VideoGLUE score (VGS), considering the tradeoff between adaptation costs and performance, to provide a holistic view of FMs’ capabilities on the video-only understanding tasks. We present the frozen-backbone evaluation results in the main paper and leave the rest to Appendix E. We employ an MAP head (Yuan et al., 2023) in action recognition (MAP probing) and spatiotemporal localization and use G-TAD (Xu et al., 2020) for temporal localization (see Section E.1 for details).
@@ 3.2 Zero-shot video-text retrieval and classification
Table 3: Results of zero-shot video-text retrieval. We report the Recall@1 (R@1) and R@5 for all the benchmarks. Note that we follow the 1K-A split of MSRVTT produced by Bain et al. (2021) which contains $1,000$ videos for testing. Please refer to Section F.3 for the results on the full split of MSRVTT proposed by Xu et al. (2016).
Methods MSRVTT (1K-A) VATEX ActivityNet
@@
VideoPrism-B 51.4 74.4 50.2 73.2 57.7 88.5 76.2 93.7 49.6 76.7 47.9 75.0
($\uparrow$7.5) ($\uparrow$4.5) ($\uparrow$4.8) ($\uparrow$4.6) ($\uparrow$4.5) ($\uparrow$5.2) ($\uparrow$2.6) ($\uparrow$0.5) ($\uparrow$6.8) ($\uparrow$7.1) ($\uparrow$7.2) ($\uparrow$7.4)
VideoPrism-g 52.7 77.2 51.7 75.2 62.5 91.0 77.1 95.6 52.7 79.4 50.3 77.1
($\uparrow$8.8) ($\uparrow$7.3) ($\uparrow$6.3) ($\uparrow$6.6) ($\uparrow$9.3) ($\uparrow$7.7) ($\uparrow$3.5) ($\uparrow$2.4) ($\uparrow$9.9) ($\uparrow$9.8) ($\uparrow$9.6) ($\uparrow$9.5)
@@
We evaluate VideoPrism’s zero-shot video-text retrieval performance on three benchmarks: MSRVTT (Xu et al., 2016; Bain et al., 2021), VATEX (Wang et al., 2019), and ActivityNet (Krishna et al., 2017). For zero-shot video classification tasks, we experiment with Kinetics-400 (Kay et al., 2017), Charades (Sigurdsson et al., 2016), SSv2-Temporal and SSv2-Events (Sevilla-Lara et al., 2021; Bagad et al., 2023), and the ATP-Hard subset of NExT-QA (Buch et al., 2022). SSv2 and NExT-QA (ATP-Hard) focus on motion and temporal reasoning, respectively. Moreover, we adapt Charades-STA (Gao et al., 2017) to the zero-shot classification scenario by reformulating each of its samples in the test set into a multi-choice retrieval problem (see Section F.2 for more details). We report results following the standard evaluation metric for each benchmark.
@@
Tables 3 and 4(e) summarize the results of video-text retrieval and video classification, respectively. VideoPrism sets the new state of the art on most benchmarks, and the gains over the prior arts are exceptionally substantial on the challenging datasets (e.g., $9.5\%$ on ActivityNet, $4.4\%$ on SSv2-Events, and $6.6$ mAP on Charades). Most results from our base-scale VideoPrism-B are actually better than those of existing larger-scale models. Additionally, VideoPrism is on par with or better than the models pretrained with in-domain data and extra modalities (e.g., audios) in Table 4(e). These improvements in zero-shot retrieval and classification tasks present VideoPrism’s strong generalization capabilities.
@@ 3.3 Zero-shot video captioning and QA
We evaluate the model in the zero-shot setting on the test splits of a suite of standard video captioning datasets including MSRVTT (Xu et al., 2016), VATEX (Wang et al., 2019), and YouCook2 (Zhou et al., 2018), and video QA benchmarks including MSRVTT-QA (Xu et al., 2017), MSVD-QA (Xu et al., 2017), and NExT-QA (Xiao et al., 2021). For video QA, where it is imperative to match the length and style of the model’s answers with groundtruths, we adopt the zero-shot approach of Flamingo (Alayrac et al., 2022) and use two-shot text-only prompts from the training set of the downstream task. Additionally, for MSRVTT-QA and MSVD-QA, we experiment with the closed-vocabulary evaluation configuration (Li et al., 2022; Yang et al., 2022). In this setting, we let the model score candidate answers according to their log-likelihoods and return the top one.
@@ 3.6 Limitations
One limitation of our approach is that we leverage a video corpus with noisy text as part of pretraining. This noisy text is potentially incomplete and biased, which could impact model performance. Moreover, long video understanding remains a challenge, since our current focus is on short video clips from which we sample $16$ frames as input to VideoPrism. Future work in this direction could leverage our encoder as part of a long video understanding system. Finally, while we advocate for the frozen-backbone evaluation, we acknowledge that there are scenarios that benefit more from end-to-end finetuning and parameter-efficient adaptation. Despite these limitations, the results demonstrate the potential impact of VideoPrism on a range of real-world video understanding tasks.
@@ 4 Related work
Foundation models (FMs) (Bommasani et al., 2021) have demonstrated tremendous promise with early work in LLMs (Devlin et al., 2019; Brown et al., 2020). Some ViFMs are built around LLMs (Wang et al., 2022d; Li et al., 2023a; Zhang et al., 2023a; Chen et al., 2023a), analyzing videos by feeding associated text to LLMs, such as ASR transcripts and machine-generated captions. In contrast, VideoPrism takes a video-centric view, and we aim to tackle a broader range of video understanding tasks.
ViFMs. Most recent FMs in CV focus on images (Radford et al., 2021; Yuan et al., 2021; Jia et al., 2021; Yu et al., 2022; Alayrac et al., 2022; Yan et al., 2022; Wang et al., 2022a; Chen et al., 2023c; Xu et al., 2023; Girdhar et al., 2023; Zhang et al., 2023b; Zhu et al., 2024). Their pretraining data contains no or only a small portion of videos, and the model architectures and learning methods are for images by design. While these FMs can accept video frames as input, they fall short on motion and temporal modeling (Yuan et al., 2023). Our work directly addresses this gap by developing a video encoder designed for video-specific applications.
@@
Pretraining strategy. Our pretraining integrates vision-language contrastive learning (Radford et al., 2021; Xu et al., 2021; Bain et al., 2022) and masked data modeling (Devlin et al., 2019; He et al., 2022). The former has led to strong late-fusion models like CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), CoCa (Yu et al., 2022), and the latter is proven effective to learn from single-modality data like language (Devlin et al., 2019; Anil et al., 2023), audio (Borsos et al., 2023), images (He et al., 2022; Wang et al., 2023d; Oquab et al., 2023), and videos (Tong et al., 2022; Wang et al., 2023b). While EVA (Fang et al., 2022, 2023) and UMT (Li et al., 2023b) transfer indirect semantics from CLIP (Radford et al., 2021) to masked modeling, we learn video-native semantics. We also introduce global distillation and token shuffling to the masked video modeling to orchestrate both appearance and motion cues.